OpenStudy (anonymous):

Find the exact value of tan(cos^-1(-3/5))

OpenStudy (ranga):



OpenStudy (ranga):

cos(x) = adjacent / hypotenuse = -3/5 (from the diagram) Therefore, cos^-1(-3/5) = x tan(cos^-1(-3/5)) = tan(x) = opposite / adjacent = 4 / (-3) = -4/3.

OpenStudy (anonymous):

Thanks!

OpenStudy (ranga):

You are welcome.